OpenStudy (anonymous):

T=2(pi)(a)(h)+2(pi)(b)(h)+R for a

OpenStudy (lgbasallote):

\[\large T = 2\pi ah + 2\pi bh + R\] first put all the non x TERMS on the left side and the x term on the right \[\large T - 2\pi bh - R = 2\pi a h\] now divide both sides by the coefficient of a (which is \(2\pi h\)) got it?

OpenStudy (anonymous):

(T-2(pi)bh-R)/2(pi)b=a is that right?

OpenStudy (lgbasallote):

close...it should be \[\large \frac{T - 2\pi bh -R}{2\pi h} = a\] you wrote 2 (pi) b for the denominator ;)

OpenStudy (anonymous):

Oooh, great, thanks